The Lager division of Suds Brewery has a current return on investment of 10%. The company has established an 8% minimum rate of return for the division.The division's manager has two investment projects available, for which the following estimates have been made: Project A Annual Income = $24,000, operating assets = $400,000 Project B Annual Income = $60,000, operating assets = $550,000 Which Project should be funded?
Answer #1
Project A ROI = 24000/400000 = 6%
Project B ROI = 60000/550000 = 10.91% or 11%
Project B should be funded